OpenStudy (anonymous):

If 9 and 12 each divide Q without remainder, which of the following must Q divide without remainder?

OpenStudy (anonymous):

A) 1 B) 3 C) 36 D) 72

OpenStudy (anonymous):

C and D

OpenStudy (anonymous):

b

OpenStudy (anonymous):

ithink its B) 3

OpenStudy (anonymous):

9 divides Q means Q is divisible by 9 Q divides X means X is divisible by Q

OpenStudy (anonymous):

\[\frac{Q}{9}=\frac{Q}{3*3}\]and \[\frac{Q}{12}=\frac{Q}{4*3}\]Thus Q must be divisible by 3, without remainder.

OpenStudy (anonymous):

It is not asking what Q is divisible by but rather what Q divides...